The angle of elevation from a point on level ground to the top of a tower is 32.34. The angle of elevation of the top of the tower from a point 122 feet further back on the same horizontal line is 22.17. Find the height of the tower to the nearest tenth of a foot.
Answer by stanbon(75887)

The angle of elevation from a point on level ground to the top of a tower is 32.34 degrees. The angle of elevation of the top of the tower from a point 122 feet further back on the same horizontal line is 22.17 degrees. Find the height of the tower to the nearest tenth of a foot.
----------
Draw the picture.
--------
tan32.34=h/x ; tan22.17=h/(x+122)
[tan32.34]x = [tan22.17)(x+122)
(tan32.34-tan22.17)x = 122[tan22.17)
0.6232329x=49.7128
x=79.77 ft.
-----------
h=x*tan32.34
h=79.77*tan32.34
h=50.50 ft. (height of the tower)
Cheers,
Stan H.